# Dynamic Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models

### v4

@@ 2.3 The introduction of DUAT
where $\mathbf{W}=\mathbf{W}_{1}\cdots\mathbf{W}_{N}$ and $\mathbf{b}=b_{1}\cdots b_{N}$igure 2 illustrates the comparison between UAT and DUAT. On the left side, Figure 2.a demonstrates the deficiency in data recognition. The input to the network is shown, with only the initial part differing slightly. This can be likened to tasks in LLMs, such as translating, continuing, or summarizing the same passage, where only the initial text segment differs while the rest remains the same. Figure 2.b provides a schematic illustration of the output for the input shown in Figure 2.a. Because the UAT can only fit a single function, the output closely follows the trend of the input, resulting in nearly identical outcomes regardless of the slight differences in the initial part. Essentially, the output remains the same regardless of the prompt.
In contrast, Figure 2.c shows the schematic output of the DUAT (Dynamic Universal Approximation Theorem) function. In DUAT, some or all parameters are functions of the input, allowing it to produce different outputs based on different prompts. This flexibility enables DUAT to generate varied results.
On the right side, Figure 2 compares UAT and DUAT through the distribution of data. Since the parameters of UAT are fixed, it can only fit a fixed data distribution. Given that we typically train with large datasets, the goal for UAT is to fit as much data as possible, which often results in fitting only the densest part of the data distribution, represented in purple. On the other hand, DUAT can dynamically adjust its parameters based on the input, allowing it to fit multiple different distributions simultaneously.
This comparison highlights the advantages of DUAT in handling more complex and diverse data scenarios, making it a more versatile and powerful tool for various applications.

### v5

@@ 2.3 The introduction of DUAT
where $\mathbf{W}=\mathbf{W}_{1}\cdots\mathbf{W}_{N}$ and $\mathbf{b}=b_{1}\cdots b_{N}$igure 2 illustrates the comparison between UAT and DUAT. On the left side, Figure 2.a illustrates the limitation of the Universal Approximation Theorem (UAT) in fitting only a single function due to its fixed parameters post-training. In contrast, Figure 2.b demonstrates the capability of the Dynamic Universal Approximation Theorem (DUAT) to approximate multiple functions. In DUAT, some or all parameters are functions of the input, allowing it to dynamically adjust the parameters that would be fixed in UAT, thereby fitting different functions based on varying inputs. This adaptability enables DUAT to generate a diverse range of outputs.
On the right side, Figure 2 compares the data distribution fitting capabilities of UAT and DUAT. Since UAT has fixed parameters, it can only fit a single, fixed data distribution. Given that training typically involves large datasets, the objective for UAT is to fit as much data as possible, which often results in fitting only the densest regions of the data distribution, highlighted in purple. Conversely, DUAT can dynamically adjust its parameters based on the input, allowing it to simultaneously fit multiple distinct distributions.
This comparison underscores the advantages of DUAT in handling more complex and diverse data scenarios, making it a more versatile tool for various applications.